Enter the equation of the following using line in slope-intercept form.
Two points on the line: (3, 0) and (5, 2)

Answer by nyc_function(2741)

Let m = slope
m = 2 - 0/5 - 3
m = 2/2
m = 1
y - y1 = m(x - x1)
y - 0 = 1(x - 3)
y = x - 3
Done!